OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 - 2x + 5

OpenStudy (anonymous):

y = a(x-h)^2+k

OpenStudy (anonymous):

@surdawi Can you help me solve it. I kind of forgot how to do it.

OpenStudy (zeesbrat3):

-b/2a then put the answer into the x values to get y

OpenStudy (anonymous):

h=-b/(2a)

OpenStudy (anonymous):

@surdawi ?????

OpenStudy (anonymous):

@zeesbrat3 what?

OpenStudy (anonymous):

y = x2 - 2x + 5 y = ax2 - bx + c

OpenStudy (anonymous):

@surdawi H=-1?

OpenStudy (anonymous):

show me how?

OpenStudy (anonymous):

@surdawi -2/2(1) =-2/2 =-1

OpenStudy (anonymous):

its -b = -(-2) =2

OpenStudy (anonymous):

@surdawi so its just 1

OpenStudy (anonymous):

yes now plug the 1 for x and solve for y y = x2 - 2x + 5

OpenStudy (anonymous):

@surdawi y=(1)^2-2(1)+5 =1-2+5 =4

OpenStudy (anonymous):

h = 1 k = 4 now plug that in to the equation i sent earlier

OpenStudy (anonymous):

@surdawi y=(x-1)^2+4

OpenStudy (anonymous):

yes